OpenStudy (raffle_snaffle):

What is the molarity of a solution of hydrochloric acid, HCL, if 5.00 mL reacts completely with 20.00 mL of NaOH to produce water and NaCl? The NaOH is known to contain 2.00 mmol of hydroxide ions (OH^-1) per milliliter. Note: mmol = millimole = 10^-10 mole.

OpenStudy (raffle_snaffle):

@ganeshie8

OpenStudy (abb0t):

This is essentially a titration problem, the addition of a strong acid to a strong base. The reaction you have happening here is basically: \(\sf \color{red}{H^+_{(aq)} + OH^-_{(aq)} \rightarrow H_2O_{(l)}}\) the others are just spectator ions since they are both strong acid/base, they completely ionize in aqueous solution. So, start by writing down what you have, 0.002 mole not sure why you have 1 mmole = 10\(\sf ^{-10}\) mole.....?

OpenStudy (abb0t):

If you have 0.002/0.02 = 0.1 M solution of NaOH that's your initial starting concentration of base.

OpenStudy (abb0t):

If you add 0.005 L of HCl, you simply add that to what you have, since it's a 1:1 ratio, you shoulldn't have a problem. So now you have: \(\sf \large \frac{0.002 ~mol}{(0.02 + 0.005) L}\) =

OpenStudy (raffle_snaffle):

what happens if you have a 1:2 ratio?

OpenStudy (abb0t):

But you don't.